# OFER: Occluded Face Expression Reconstruction

### v1

@@ 8.2 Identity Ranking Network
Table 5: Ablation Study on the Importance of Ranking. Ideal Lowest Error Sample refers to the sample with the lowest median MSE compared to the ground truth scan. Ranked Sample is the optimal sample selected by the Ranking Network from all samples generated by the Identity Generative Network. Average of All Samples represents the average error across all generated samples. ”NA” in the Ranked Sample column for the 200-1K range indicates that our network, which was trained on 100 samples, cannot validate these larger sample sizes.
Quantity of input samples.
@@
As detailed in Sec. 4.2, increasing the number of generated samples results in a more thorough coverage of the data distribution, such that at least one sample– the optimal sample–is closer to the ground truth. Therefore, as sample set size increases and leads to denser mode coverage, the optimal sample is expected to get closer to the ground truth, resulting in a decrease in error between the optimal sample and the ground truth. However, for the ranking network to identify that optimal element, it must be be trained on a sufficiently large number of samples to effectively rank them and select the best one. This requirement comes with a significant computational costs, including increased time, especially due to the slower inference times of diffusion models (ref. Sec. 4.2). Thus, finding the right balance in the number of samples used for training is important. We conducted an ablation study on different sample sizes generated by the Identity Generative Network and assessed the median error of the optimal sample from each set to determine the ideal number of samples for training the Identity Ranking Network. The results are presented in  Tab. 6.
#samples Optimal sample
@@
Table 6: Ablation on effect of number of samples in selection of optimal sample which gives low median MSE on reconstructed samples. The results are evaluated on NoW validaion benchmark
Input to Ranking Network.
Selecting the most accurate sample requires a network design that excludes irrelevant information (See Sec. 4.2 for more details). To validate this claim, we conducted an ablation study comparing three inputs: (a) shape coefficients, which are a PCA model representing entire head shape; (b) all landmark vertices of the reconstructed head shape; and (c) front-face vertices, $x_{ssf}$, from the reconstructed head shape. The results of this comparison are presented in Tab. 7. The high errors associated with the coefficients and vertex-based reconstructions likely stem from the inclusion of irrelevant information.
@@
The challenge in selecting the best representative shape from the samples generated by the Identity Generative Network is that, even with mild occlusions where most of the facial structure is preserved, variability can still occur. This means we need an appropriate loss function that mitigates the selection of sub-optimal samples. We considered two loss functions for this network: binary cross-entropy and softmax loss, which approximates the ground truth error distribution considering all samples. To assess the performance of both, we conducted an ablation study, the results of which are presented in  Tab. 8. Since more than one sample can be optimal, ranking and selection using softmax loss yields better precision with a set of higher-ranked samples.
@@
(a) Shape coefficients($\hat{S}$) 3.3 21.3 33.2 26.7 40.0 53.3 0.54 0.57 0.64
(b) Vertex($x_{ssf}$) 3.3 11.3 24.8 13.3 30.0 53.3 0.49 0.58 0.61
(c) MeanResidual ($X_{\text{mean}},x_{\text{res}}$) 3.3 44.5 64.7 33.3 80.0 93.3 0.68 0.78 0.83
Table 7: Ablation for input to ranking network. (a) is the 300 shape coefficients generated from Identity Generative Network; (b) is the front face vertices $x_{\text{ssf}}$, 1785 out of the 5023 ($x_{ss}$), of FLAME from the reconstructed head shape using coefficients of (a); (c) is the concatenated input of mean($X_{\text{mean}}$) of (b) and the residual vertices ($x_{\text{res}}$) obtained by subtracting $x_{\text{ssf}}$ from $X_{\text{mean}}$.
Loss precision IOU error (GT/Pred)
@@
Table 8: Ablation for loss function for Identity Ranking Network. We trained the network with Binary Cross Entropy (BCE) loss and Cross Entropy on Softmax (Softmax) loss using 100 Stirling(HQ) [23] frontal face images. For validation, we used 20 Florence [3] dataset frontal face images. IOU represents intersection over union of predicted ranking order and ground truth ranking order for first 10, 20 and 30 sorted rank index. error(GT/Pred) shows the average error for first 1, 5 and 10 ground truth rank samples and that of predicted rank samples.
@@ 9.1 Training Dataset
In Tab. 9, we list the datasets used to train the networks in our framework. FaMoS [6] dataset comprises of 3D registered FLAME meshes. We utilized only a subset of them such that it covered all the expression variations in the entire dataset. Please note that our network trains in parametric space rather than on meshes. Therefore, we obtained the corresponding parameters for this subset directly from the authors of FaMoS. We used the expression coefficients from this subset to train the Expression Generative Network. The remaining four datasets listed in the table (Stirling [23], Florence [3], Facewarehouse [8], LYHM [9]) were used to train Identity Generative Network and Identity Ranking Network.
@@
Table 9: Datasets used for training Networks. FaMoS expression coefficients are used to train Expression Generative Network. The shape coefficients from the remaining four datasets were used to train Identity Generative Network and Identity Ranking Network.
@@ 10.1 Quantitative Results
Method Average error of samples
Med $\downarrow$ Mean $\downarrow$ std $\downarrow$
FLAME(Baseline) Without ranking 1.79 2.25 1.90
DECA [22] 1.17 1.46 1.25
MICA (4DS) [55] 1.02 1.25 1.05
MICA (8DS) [55] 0.90 1.11 0.92
TokenFace [52] 0.79 0.99 0.85
FOCUS [31] 1.05 1.31 1.14
FOCUS (MP) [31] 1.03 1.29 1.12
Diverse3DFace[14] 1.41 1.78 1.52
OFER (ours) (4DS) 1.02 1.25 1.04
OFER (sample selected by ranking) 0.98 1.21 1.01
Table 10: Comprehensive comparison of non-metrical evaluation in NOW validation benchmark. FLAME (baseline) is created by sampling random FLAME shape coefficients. “DS” stands for Dataset. We obtained only 4 out of the 8 datasets used by MICA. We evaluate using the median (Med), mean, and standard deviation (std) of the non-metrical Mean Square Error (MSE) between the reconstructed shapes and ground truth scans. The last row shows our results using the sample selected by the ranking network.
@@
Table 11: Neutral face 3D Metrical reconstruction error on the NoW test benchmark.. The results show comparison of accuracy of single view reconstruction methods based on NoW benchmark challenge.
@@
Figure 13: Comparison of expression reconstruction for in-the-wild occluded images. We compare against EMOCA [10] (showing front and side view, pink), two reconstructions from Diverse3D [14] (blue), and three samples (front and side view) from our method (green).
Figure 14: Ranking on NoW validation. Column 1 shows the optimal sample selected by Identity Ranking Network, column 2 displays the sample ranked 5th, and the last column shows the lowest-ranked sample. Although error differences are subtle, variations can be observed between the higher-ranked samples (rank 1 and 5) and the lower ranked-sample (rank 100) in the nose and lip regions of image (a), the eye region of image (b), and the lip and chin region of image(c). This demonstrated the effectiveness of ranking in selecting higher-quality samples.

### v2

@@ Results.
Table 1: Ablation study on the importance of ranking. Ideal Lowest Error Sample refers to the sample with the lowest median MSE. Ranked Sample is the optimal sample selected by the Ranking Network from all samples generated by the IdGen. Average of All Samples represents the average error across all generated samples. “NA” in the Ranked Sample column for the 200-1K range indicates that our network, which was trained on 100 samples, cannot validate these larger sample sizes.
Method Unoccluded Occluded Both
Med Mean std Med Mean std Med Mean std
FLAME 1.79 2.24 1.89 1.80 2.26 1.91 1.79 2.25 1.90
Deep3D [12] 1.33 1.67 1.41 1.40 1.73 1.41 1.36 1.70 1.41
DECA [22] 1.18 1.47 1.24 1.29 1.56 1.29 1.17 1.46 1.25
MICA (4DS) [59] n/a n/a n/a n/a n/a n/a 1.02 1.25 1.05
MICA (8DS) [59] n/a n/a n/a n/a n/a n/a 0.90 1.11 0.92
TokenFace [56] n/a n/a n/a n/a n/a n/a 0.79 0.99 0.85
FOCUS [31] 1.03 1.25 1.03 1.07 1.34 1.19 1.05 1.31 1.14
FOCUS (MP) [31] 1.02 1.24 1.02 1.08 1.34 1.20 1.03 1.29 1.12
Diverse3D [13] n/a n/a n/a n/a n/a n/a 1.41 1.78 1.52
OFER-rank (ours) 0.97 1.20 1.00 1.01 1.26 1.05 0.98 1.21 1.01
OFER-min 0.81 1.04 0.91 0.84 1.10 0.96 0.81 1.05 0.92
OFER-avg 1.01 1.25 1.03 1.04 1.29 1.08 1.01 1.25 1.04
Table 2: Quantitative evaluation on the NoW validation benchmark, on the unoccluded, occluded and full subsets. “FLAME” is created by sampling random FLAME shape coefficients. Rows 2-6 show standard reconstruction methods, while the bottom rows show occlusion-based approaches. OFER-rank is our result using IdGen and IdRank networks. The bottom two rows show the minimum (OFER-min) and average (OFER-avg) reconstruction errors obtained over the 100 samples provided by IdGen. We show median (med), mean, and standard deviation (std) of the non-metrical MSE between the reconstructed and ground truth shapes.
Figure 8: Qualitative and quantitative results for best and least ranked samples for row (b) in Tab. 1 evaluated on NoW validation images. The median MSE error between the ground truth scan and the reconstructed 3D shape is displayed below each rendering. These results demonstrate that ranking as a selection method enhances the quality of sample selection.
4.3 Evaluation on reconstructions with ExpGen
We run experiments on the full expressive reconstructions enabled by ExpGen. Similarly to IdGen, the network was trained for 400 epochs with a batch size of 128 on an RTX 8000 GPU, with the training spanning one GPU day.
Figure 9: Comparison of expression reconstruction for in-the-wild occluded images. We compare against EMOCA [9] (showing front and side view, pink), two reconstructions from Diverse3D [13] (blue), and six samples (front and side view) from our method (green).
Method $N_{C}^{unocc}$ $N_{C}^{occ}$ $E_{D}^{occ}$ $E_{C}^{unocc}$ $E_{C}^{occ}$
Diverse3D[13] 2.44 2.16 6.4 7.46 4.72
OFER (ours) 1.90 1.95 3.2 3.48 2.78
Table 3: Quantitative evaluation using SE on unoccluded ($unocc$) and occluded images ($occ$) in Dey and CO-545 datasets. The letters represent: $N$ for Neutral, $E$ for expression, $D$ for Dey dataset, and $C$ for CO-545 dataset. Reconstructions evaluated on all front face vertices for $unocc$ images and only on unoccluded vertices for $occ$ images.
Method CSE minASD-V maxASD-V ASD-O ODE ODE-O ODE-V
RMSE $\downarrow$ RMSE $\downarrow$ RMSE $\downarrow$ HM $\uparrow$ max md $\downarrow$ max md $\downarrow$ max md $\downarrow$
Diverse3D[13] 0.30 1.31 1.96 0.16 0.4 0.57 0.33
OFER (ours) 0.17 0.65 1.39 0.18 0.01 0.01 0.01
Table 4: Quantitative evaluation of expression reconstructions in terms of CSE, ASD-V and ODE evaluated using CO-545 dataset. The metrics are details in  Sec. 4.3
Method STD-S ODE
$mask$ $sunglasses$ $naturalocc$ $mask$ $sunglasses$ $naturalocc$
RMSE $\uparrow$ RMSE $\uparrow$ RMSE $\uparrow$ RMSE $\downarrow$ RMSE $\downarrow$ RMSE $\downarrow$
Diverse3D[13] 11.81 21.28 20.84 0.95 0.34 0.41
OFER (ours) 34.04 34.38 34.56 0.002 0.001 0.003
Table 5: Quantitative evaluation of expression reconstructions using STD-S and ODE metrics evaluated on 200 images per subset (mask, sunglasses and natural occlusions) from  [20].
Baselines and Metrics.
We compare our method against the state-of-the-art diverse expression reconstruction approach by Dey et al. [13] and the single-hypothesis expression reconstruction method EMOCA [9]. It is important to note that Dey et al. [13] employs test-time optimization, whereas our approach does not. Additionally, their network is trained on the COMA dataset, the same dataset used in the creation of CO-545. As a result, the evaluation data overlaps with their training set, potentially giving their results an additional advantage.
Building on the metrics introduced by  [54] and  [13] we evaluate on six metrics: (1) Sample Error (SE): the average Root Mean Square Error (RMSE) between the landmark vertices of the reconstructed samples and the ground truth; (2) Closest Sample Error (CSE): the SE of the closest reconstructed sample; (3) Average Self Distance-Visible (ASD-V): to ensure that the visible regions of the reconstructions remain close to the input, the maximum distance across samples should be minimized. maxASD-V measures the average per-vertex RMSE on the visible regions of a sample and its “farthest” neighboring sample across all instances, while minASD-V measures ASD-V with the “nearest” neighbour sample; (4) Average Self Distance-Occluded (ASD-O): the reconstructions of occluded regions should show high diversity, given by $d^{o}=\frac{\text{RMSE(gt}^{v}{\text{,x}_{i}^{v}}{\text{)}}}{\text{max(MD)}^{o}}$, while the visible regions must remain close to the ground truth, given by $d^{v}=\frac{\text{RMSE(x}_{i}^{o}{\text{,x}_{j}^{o}}{\text{)}}}{\sqrt{\text{max(MD)}^{o}}}$. Here, $\text{max}(MD^{o})$ is the maximum mahalanobis distance threshold of the occluded vertices, and $x_{j}$ is the nearest neighbour sample of $x_{i}$ based on the L2 distance of occluded vertices. The superscript $v$ and $o$ point to the visible and occluded vertices. We compute ASD-O as the harmonic mean (HM) between $d^{v}$ and $d^{o}$, given by  Eq. 7; (5) Shape Standard Deviation (STD-S): in case of absence of ground-truth 3D data, we opt for the standard deviation metric of the entire shape, excluding data which falls outside the maximum mahalanobis distance per vertex calculated using COMA [42]; (6) Out-of-Distribution Error (ODE): ODE measures the average per-vertex mahalanobis distance (MD) out-of-distribution error. It addresses the observation that reconstructions from Dey et al. [13] occasionally produce non-plausible expressions that do not fall within the distribution of ground truth expressions.
[…]
Results.
The qualitative comparison of diverse expressive faces for occluded images is shown in Fig. 2 and Fig. 9. Our approach effectively generates plausible faces with varied expressions, even under challenging occlusions such as face masks. In contrast, EMOCA [9] tends to produce degenerate faces in similar hard occlusion scenarios, while Dey et al. [13] often results in unrealistic and extreme reconstructions. The quantitative evaluation shown in Tab. 3,  Tab. 4 and  Tab. 5 also supports this, particularly highlighting the substantial improvement over Dey et al. [13]. We sampled 15 expressions per image for this evaluation. Our method demonstrates a significant advantage, reliably reconstructing visible regions and producing varied expressions in occluded areas.
5 Limitations and Future work
The generative capabilities of diffusion models with diverse inputs are attributed to the amount and variations in the training data [46]. In our case, we had access to only a small subset of the available 3D supervision dataset for both IdGen and ExpGen, As a result, our generated expressions may lack certain expressive details. Therefore, a promising direction for future research is to integrate both 2D and 3D supervision to improve the performance of the method.
When we refer to performance, we specifically measure it in terms of how closely the reconstructed model approximates the ground-truth scan. Although our ranking network selects a sample that minimizes the error compared to a randomly chosen one from the generated set, it does not guarantee the selection of the absolute best match, i.e., the sample with the lowest error and closest approximation to the ground truth. From this observation we identify two potential research directions for constructing an ideal ranking network. The first is to refine the ranking architecture itself by adopting a more suitable loss function. Softmax scoring is not ideal when presented with large number of sample sets since it dilutes the scores, making it challenging to identify the best among the higher-quality samples. Additionally a more sophisticated ranking mechanism [6] can complement to the selection of quality samples. A second promising approach involves integrating the ranking process as a feedback mechanism during training of the diffusion model. Rather than using a stand-alone model, this integration could improve the quality of the generated samples, resulting in improved overall performance.
6 Conclusion
In this paper we introduced OFER, a conditional diffusion-based method for generating multiple hypotheses of expressive faces from a single-view, in-the-wild occluded image. Key to OFER is the use of two diffusion models to generate FLAME 3DMM shape and expression coefficients. To ensure a consistent geometric face shape for varied expression of a single identity, we introduced a probabilistic ranking method to select an optimal sample from the generated shape coefficients. By combining the statistical learning strengths and generative capabilities of diffusion models, along with the smooth face reconstruction provided by a parametric model, our method produces plausible 3D faces that accurately reflect the input image. OFER achieves state-of-the-art results in diverse expression reconstruction outperforming existing occlusion-based methods, and can generate plausible and diverse results for a given input.
Supplementary Material
7 Ablation Study
7.1 Identity Generative Network (IdGen)
Choice of image embedding.
The representation used for the conditioning image provided to the diffusion network plays an important role in the reconstruction task. In Tab. 6 we show ablation results for the embedding used in IdGen. The results indicate that using only the ArcFace [10] embedding improves the error metric of this network, compared to employing FaRL [57] or a combination of both as conditions. This may be because ArcFace [10] is used to distinguish identity-defining features, while FaRL [57] is designed for downstream facial analysis tasks that require capturing facial subtleties. In addition, we compare the performance of ArcFace with state of the art image embeddings, DINOv2 [36] and CLIP [40]. The results are presented in Tab. 7. The reason we hypothesize for the lower performance of CLIP and DINOv2 is that they are trained on multi-domain models and not specifically trained on faces like ArcFace.
7.2 Identity Ranking Network (IdRank)
Ranking selects low-error samples
We showcase the ability of ranking network to identify subtle differences to select top-ranked sample with lower MSE error than the lower ranked samples in Fig. 10.
Embedding Occluded subset Unoccluded subset
ArcFace [10] 1.03 1.26 1.06 1.01 1.22 1.01
ArcFace+FaRL 1.05 1.30 1.09 1.03 1.28 1.06
FaRL [57] 1.15 1.40 1.12 1.07 1.30 1.06
Table 6: Ablation on image embeddings for the IdGen network. MSE error on the NoW validation dataset, using ArcFace [10] embedding, FaRL [57] embedding, or a combination of both.
Figure 10: Ranking on NoW validation images. For the image in column (a), we generate 100 shape coefficients with IdGen and rank them using IdRank. Column (b) shows the top-ranked sample, (c) a mid-ranked sample (50th), and (d) the worst-ranked sample. Columns (b)-(d) show viewpoints highlighting the sample errors, with its median MSE shown underneath each sample.
Method Arcface CLIP DINOv2
Median 1.01 1.65 1.68
Mean 1.24 2.07 2.08
Table 7: Ablation study for IdGen embedding on NoW validation benchmark. ArcFace trained for identity clustering performes better in comparison to other image embedding for identity recognition reconstruction task.
Sampling Quantity for Input.
@@ 7.2 Identity Ranking Network (IdRank)
Naturally, increasing the number of generated samples results in a more thorough coverage of the data distribution. As the sample set size increases and leads to denser mode coverage, we expect the average error to get closer to the ground truth. However, for the ranking network to identify that optimal element, it must be be trained on a sufficiently large number of samples to effectively rank them and select the best one. This requirement comes with a significant computational costs, especially due to the slower inference time of diffusion models. Thus, finding the right balance in the number of samples used for training is important. We conducted an ablation study on different sample sizes generated by IdGen, and assessed the median error of the optimal sample from each set to determine the ideal number of samples for training IdRank. The results are presented in  Tab. 8.
#samples Ideal lowest error sample
@@
Table 8: Ablation on effect of sampling quantity as input to IdRank. To evaluate the tradeoff between the number of generated samples by IdGen and the ”ideal” lowest error sample in the set. The values are the lowest median error sample evaluated against the Now validation benchmark.
Face vertices as Input.
Selecting the most accurate sample requires a network design that excludes irrelevant information (See Sec. 3.3 for more details). To validate this claim, we conducted an ablation study comparing three inputs: (a) shape coefficients, which are a PCA model representing entire head shape; (b) all landmark vertices of the reconstructed head shape; and (c) front-face vertices from the reconstructed head shape. The results of this comparison are presented in Tab. 9. The high errors associated with the coefficients and vertex-based reconstructions likely stem from the inclusion of irrelevant information.
@@
The challenge in selecting the best representative shape from the samples generated by the IdGen is that, even with mild occlusions where most of the facial structure is preserved, variability can still occur. This means we need an appropriate loss function that mitigates the selection of sub-optimal samples. We considered two loss functions for this network: binary cross-entropy and softmax loss, which approximates the ground truth error distribution considering all samples. To assess the performance of both, we conducted an ablation study, the results of which are presented in  Tab. 10. Since more than one sample can be optimal, ranking and selection using softmax loss yields better precision with a set of higher-ranked samples.
@@
(a) $\hat{S}$ 3.3 21.3 33.2 26.7 40.0 53.3 0.54 0.57 0.64
(b) $M^{\text{frontal}}$ 3.3 11.3 24.8 13.3 30.0 53.3 0.49 0.58 0.61
(c) X ($\mu_{M},M^{\prime}$) 3.3 44.5 64.7 33.3 80.0 93.3 0.68 0.78 0.83
Table 9: Ablation for input to ranking network. $\hat{S}$ is the 300-dimentional shape coefficients generated from IdGen; $M^{\text{frontal}}$ is the front face vertices of FLAME mesh $M$ reconstructed from ($\hat{S}$ ; X=($\mu_{M},M^{\prime}$) is the mean, residual pair defined in  Sec. 3.3
[…]
@@
Table 10: Ablation for loss function for IdRank. We trained the network with Binary Cross Entropy (BCE) loss and Cross Entropy on Softmax (Softmax) loss using 100 Stirling(HQ) [23] frontal face images. For validation, we used 20 Florence [3] dataset frontal face images. IOU represents intersection over union of predicted ranking order and ground truth ranking order for first 10, 20 and 30 sorted rank index. error(GT/Pred) shows the average error for first 1, 5 and 10 ground truth rank samples and that of predicted rank samples.
@@
8 Design Choices
8.1 Ranking by distribution matching.
The objective of the ranking model is to sort the output from the network and optimize a loss based on ranking order. We choose a list-wise ranking method due to its demonstrated effectiveness [6]. Since sorting and ranking are non-differentiable operations, we reformulate the sort-rank problem into a probability distribution alignment problem aiming to minimize the softmax loss $L_{R}$ between the ground truth distribution $g$ and the predicted distribution $h$.
8.2 Exclusion of Ranking in Expression Generation
Our ranking network is trained to rank only neutral shapes and not expressions. This is because shape geometry remains consistent despite occlusions or variations in expressions in the input image. But, ranking the expression hypotheses is harder since multiple hypotheses can be equally valid for occluded regions.
9 Architecture
The detailed overview of our 1-D Unet-transformer hybrid architecture of our IdGen and ExpGen networks is shown in Fig. 12. Both share similar architecture, differing in the embeddings and the inputs. The conditional embedding of IdGen obtained from ArcFace [10] is $\mathbb{R}^{512}$, and the embedding of ExpGen obtained by concatenating ArcFace [10] and FaRL [57] embeddings is $\mathbb{R}^{1024}$.
Figure 12: 1D Unet Transformer Hybrid architecture of IdGen and ExpGen. Each block of the UNet is comprised of two resnet blocks, followed by a self-attention module and a downsampling/upsampling module for the encoder and decoder respectively. Each resnet block consists of two 1D convnet followed by SiLU activation and residual connection. For IdGen, the conditional embedding ($\mathbb{R}^{512}$) is passed to every layer of the UNet. The training input to IdGen is FLAME [32] shape coefficient, $S\in\mathbb{R}^{300}$ which gets downsampled to $50$ and to $10$ at the bottleneck. For ExpGen, the conditional embedding is ($\mathbb{R}^{1024}$) and the input is FLAME expression (including 3 jaw cofficients) $E\in\mathbb{R}^{53}$ which is downsampled to $25$ and to $10$ at the bottleneck.
@@ 10.1 Training Dataset
In Tab. 11, we list the datasets used to train the networks in our framework. FaMoS [5] dataset comprises of 3D registered FLAME meshes. We utilized only a subset of them such that it covered all the expression variations in the entire dataset. Please note that our network trains in parametric space rather than on meshes. Therefore, we obtained the corresponding parameters for this subset directly from the authors of FaMoS. We used the expression coefficients from this subset to train the ExpGen. The remaining four datasets listed in the table (Stirling [23], Florence [3], Facewarehouse [7], LYHM [8]) were used to train IdGen and IdRank.
@@
Table 11: Datasets used for training Networks. FaMoS expression coefficients are used to train ExpGen. The shape coefficients from the remaining four datasets were used to train IdGen and IdRank.
@@ 11.1 Quantitative Results
Table 12: Neutral face 3D Metrical reconstruction error on the NoW test benchmark.. The results show comparison of accuracy of single view reconstruction methods based on NoW challenge.
@@
Figure 14: Comparison of expression reconstruction for in-the-wild occluded images. We compare against EMOCA [9] (showing front and side view, pink), two reconstructions from Diverse3D [13] (blue), and six samples (front and side view) from our method (green).
Figure 15: Ranking on NoW validation. Column 1 shows the optimal sample selected by IdRank, column 2 displays the sample ranked 5th, and the last column shows the lowest-ranked sample. Although error differences are subtle, variations can be observed between the higher-ranked samples (rank 1 and 5) and the lower ranked-sample (rank 100) in the nose and lip regions of image (a), the eye region of image (b), and the lip and chin region of image(c). This demonstrated the effectiveness of ranking in selecting higher-quality samples.